Grade 12th passIIT JEE Entrance Exam

If A and B are the points (-3,4) and (2,1) then the coordinate of the point C on AB such that AB -2BC are ???

asifalishah
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
If
AB = 2BC
Then
On AB, C can lie inside or outside.
If C is out side then
B is inside such that
AB/BC = 2/1
Now
Let coordinates of C = (x,y)
2 = 2x -3 /3
2x = 9
x = 9/2
And
1 = 2y +4/3
2y = -1
y = -1/2
 
Regards
Arun